# Enhancing Edge Detection by Texture Handling Architecture and Noiseless Training Data

### v4

@@ Abstract
Image edge detection (ED) is a fundamental task in computer vision. While convolution-based models have significantly advanced ED performance, achieving high precision under strict error tolerance constraints remains challenging. Furthermore, the reliance on noisy, human-annotated training data limits model performance, even when the inputs are edge maps themselves. In this paper, we address these challenges in two key aspects. First, we propose a novel ED model incorporating Cascaded Skipping Density Blocks (CSDB) to enhance precision and robustness. Our model achieves state-of-the-art (SOTA) performance across multiple datasets, with substantial improvements in average precision (AP), as demonstrated by extensive experiments. Second, we introduce a novel data augmentation strategy that enables the integration of noiseless annotations during training, improving model performance, particularly when processing edge maps directly. Our findings contribute to a more precise ED architecture and the first method for integrating noiseless training data into ED tasks, offering potential directions for improving ED models. Codes can be found on https://github.com/Hao-B-Shu/SDPED.
@@
Edge Detection, Convolutional Neural Network, Data Augmentation, Noiseless Data, Evaluation Benchmark.
@@
organization=Shenzhen University,city=Shenzhen, postcode=518060, state=Guangdong, country=P. R. China
@@
An edge detection model based on CSDB.
A data augmentation method allowing noiseless data.
A discussion of the choice of evaluation standard.
Significant improvements over high standard evaluation.
Extensive experiments implementation.
@@ 1 Introduction
Image edge detection (ED) is a fundamental task in computer vision and has been studied for decades, with extensive applications in high-level tasks such as object detection[1] and image segmentation[2]. Early ED methods, such as the Sobel operator[3] and the Canny algorithm[4], relied on gradient-based features. Later, more sophisticated statistical approaches were introduced[5; 6]. In recent years, machine learning has become indispensable in nearly all computer vision tasks, leading to the emergence of numerous learning-based ED algorithms[7; 8; 9; 10; 11; 12]. Currently, convolutional neural network (CNN)-based models achieve state-of-the-art (SOTA) performance in ED tasks [13; 14; 15; 16; 17; 18; 19], while recently, transformer-based approaches have also been explored[20].
The goal of ED is to identify edges in an image, which typically correspond to pixels with significant changes in intensity compared with their neighbors. However, textural areas can also contain rapid intensity changes that do not correspond to edges. Classical ED algorithms such as Canny[4], despite their efficiency, often fail to differentiate between edges and textures due to their inability to adapt to varying data distributions. Consequently, learning-based models, particularly CNNs, have become the dominant approach. Over the past decade, CNN-based ED models have demonstrated remarkable progress, with models such as HED[14], RCF[21], RINDNet[22], BDCN[23], DexiNed[19], and even lightweight models such as TEED[24] achieving $F$-scores comparable to or exceeding human performance in standard evaluation procedures.
@@
    Limited precision of existing models: Current ED models achieve high scores only when evaluated with a large error tolerance. For example, the standard error tolerance radius is approximately 4.3 pixels for BSDS500[6], 11.1 pixels for BIPED[25], and even 16.3 pixels for MDBD[26]. However, when stricter evaluation criteria are applied, the performance of these models drops significantly. Hence, there is a pressing need for more precise ED models that perform well under stricter evaluation settings.
@@
    Suboptimal predictions on noiseless inputs: Existing models often produce suboptimal edge maps even when given noiseless input data, namely ground-truths themselves [^1] In this paper, we define noiseless data as ground-truth edge maps, while noisy data refers to standard images containing textures. Models should ideally perform better on noiseless inputs than on noisy ones., as shown in Fig 1. Ideally, an ED model should yield nearly perfect predictions on noiseless inputs since they are already well-defined edges. Yet, most models generate results on noiseless inputs that are similar to predictions on textured images. Addressing this issue is critical for improving ED model reliability.
Figure 1:  Predictions of three previous ED models on the BRIND dataset. $Off_{E}$ denotes models retrained using the official data augmentation and tested on edge maps, while $Our_{E}$ represents models retrained using our proposed data augmentation (detailed in Section III) and tested on edge maps. Similarly, $Off_{O}$ and $Our_{O}$ correspond to models trained with respective augmentation strategies but tested on standard images. Our augmentation method produces sharper predictions on edge maps and potentially enhances performance on standard images.
To address these challenges, we propose a novel ED model incorporating cascaded skipping density blocks (CSDB). Density blocks (DBs) have demonstrated effectiveness in tasks such as image super-resolution [27], however, they perform suboptimally when directly applied to ED. Therefore, we introduce CSDB, a structural refinement tailored for ED tasks, and design new ED models based on this architecture. Additionally, to improve model performance on noiseless data, we introduce a novel data augmentation strategy that incorporates noiseless samples. Unlike conventional augmentation methods that rely solely on textured images with noisy labels, our approach allows models to learn directly from noiseless data, thereby mitigating the issue discussed in 2) and Fig. 1.
@@
    Development of high-precision ED models: We introduce a novel ED model, Skipping Density Precise Edge Detection (SDPED), which achieves SOTA performance with fewer parameters. SDPED employs deep architectures with CSDB, a structure previously unexplored in ED. Notably, our models do not require multi-scale supervision, challenging the widely held belief that such supervision is essential for ED.
@@
    Introduction of a novel data augmentation strategy: We propose a data augmentation method that incorporates noiseless edge maps into training, enabling models to achieve improved performance on noiseless data. This method provides a potential solution to the issue described in 2) and Fig. 1.
@@
    Comprehensive experimental validation: We conduct extensive experiments to validate our approaches, demonstrating their effectiveness through quantitative and qualitative analyses.
@@ 2.1 Datasets, Label Refinement, and Data Augmentation
Early ED datasets were primarily derived from segmentation tasks, such as BSDS300 and its later extension, BSDS500[28], which remains one of the most widely used benchmarks. BSDS500 consists of 500 RGB images annotated by multiple users and is frequently utilized for both segmentation and boundary detection tasks. Another well-known dataset is NYUD/NYUD2[29], which contains hundreds of thousands of RGB-D images, originally designed for indoor scene analysis. Among these, 1,449 images are annotated with segmentation categories. Other commonly used datasets include PASCAL VOC[30], Microsoft COCO[31], SceneParse150[32], and Cityscape[33]. However, since these datasets were not explicitly designed for ED, their labels may lack precision in edge annotations.
With the increasing focus on ED tasks, several dedicated datasets have been proposed. In 2016, the MDBD dataset[26] was introduced, containing 100 high-resolution (1280$\times$720) images, each annotated by multiple annotators. While initially designed for boundary detection, it has been widely adopted in ED research. In 2020, the BIPED dataset[25] was specifically developed for ED tasks, featuring 250 high-resolution (1280$\times$720) RGB images with precise edge annotations, later refined in its BIPED2 version. In 2021, BRIND[22] was introduced as a re-annotation of BSDS500, categorizing edges into four types: reflectance, illumination, normal, and depth edges. More recently, UDED[24], proposed in 2023, consists of 29 high-quality images with carefully labeled edges, curated from multiple datasets.
A fundamental challenge in all ED datasets is annotation noise. Since edge labels are based on human annotations, perfect accuracy is unattainable, and inconsistencies arise both between annotators and even within the same annotator across multiple attempts. As a result, various studies have focused on refining human-labeled data or enhancing model robustness to noisy annotations[34; 35].
Regardless of the dataset used, data augmentation is widely recognized as essential in ED tasks. Common augmentation techniques include rotation, flipping, cropping, resizing, and gamma transformations. However, augmentation strategies vary across studies. For instance, in HED[14], images are rotated at 16 different angles, flipped, and cropped to obtain 32 augmented images per sample. In RIND[22], only four rotations with flips (without cropping) are applied, resulting in eight augmented images per sample. In contrast, Dexi[19] employs a more aggressive augmentation strategy, incorporating 16 rotations, flips, cropping, and three gamma transformations, generating 288 augmented images per sample. While extensive augmentation increases data diversity, its impact on model performance remains an open question, but excessive augmentation can significantly increase training costs.
@@ 2.2 CNN-based ED Models and Loss Functions
Currently, CNN-based methods dominate ED tasks. The HED model[14] introduced a multi-scale and multi-level output strategy, where predictions at different levels are fused, and the loss is computed as the average of all supervised outputs. RCF[21] extends this approach by integrating multi-scale outputs with feature fusion across different layers, generating side outputs to enhance edge representation. BDCN[23] applies layer-specific supervision, further improving ED accuracy. PiDiNet[17] introduces pixel-difference convolutions, while RIND[22] predicts four edge types through a three-stage processing pipeline. Dexi[19] leverages extremely dense skip connections to enhance feature propagation.
Most ED models struggle to produce crisp edges without additional post-processing techniques such as non-maximal suppression (NMS). This issue may arise from the use of down-sampling operations such as max-pooling, as well as the widespread use of weighted binary cross-entropy (WBCE) as the loss function. Various strategies have been proposed to address this challenge. For instance, [36] suggests improving edge sharpness by aggregating features at larger scales during up-sampling, while [37] explores the effectiveness of Dice loss in ED tasks. More recently, tracing loss[18] has emerged as a promising alternative and has been adopted in models such as TEED[24]. However, these solutions often introduce trade-offs, requiring careful evaluation in different scenarios.
@@ 2.3 Evaluation Metrics for ED
The standard evaluation framework for ED models includes three key metrics: Optimal Dataset Score (ODS), Optimal Image Score (OIS), and Average Precision (AP). These metrics are derived from the $F_{\beta}$-score, defined as:
@@
Most ED models utilize the evaluation algorithm from[7], which includes a hyperparameter, maxDist, representing the maximum allowable error tolerance distance. The error tolerance distance is expressed as a fraction of the diagonal length of the image. For example, in BSDS500, where images have a resolution of 321$\times$481 pixels, setting $\textit{maxDist}=0.0075$ results in an error tolerance of approximately 4.3 pixels, computed as $\sqrt{481^{2}+321^{2}}\times 0.0075$. Notably, the choice of maxDist is often inconsistent across datasets. For instance, in BSDS500[28], maxDist is set to 0.0075 (about 4.3 pixels), whereas in BIPED[19], the same value translates to 11.1 pixels due to higher image resolution. In MDBD[26], maxDist is often increased to 0.011, allowing a tolerance of approximately 16.3 pixels. A stricter evaluation standard would involve reducing the error tolerance distance, though such adjustments require additional discussions.
@@ 3 Methodology
This section presents the core methodology of our approaches, including the architecture of the proposed SDPED model and an improved data augmentation strategy designed to enhance model performance on noiseless data. First, we introduce the SDPED model, which leverages the CSDB structure to improve edge feature extraction. Then, we discuss our modified data augmentation method, which incorporates noiseless data into training to mitigate the limitations caused by noisy human annotations.
3.1 The SDPED Model
The performance of ED models largely depends on their ability to suppress textural regions while preserving true edges. Since textures lack a strict definition, designing theoretically sound architectures to handle textures remains a challenge. However, ED research has progressed more slowly compared to other computer vision tasks, such as image super-resolution (SR), and the latter has produced numerous architectures experimentally validated for texture processing. Borrowing structural designs from related tasks may, therefore, be a promising direction for improving ED models.
Based on this observation, we propose the SDPED model, illustrated in Fig. 2, with additional visualizations in Fig. LABEL:fig:Legend and 3. The model employs CSDBs in the main processing path, generating multiple outputs that are subsequently fused by a dedicated fusing block. The Dense Block (DB) structure has been widely used in SR tasks [38; 27], demonstrating strong performance in texture processing. Since texture-rich regions pose challenges in both SR and ED tasks, it is reasonable to explore similar architectures for ED. However, to better suit ED-specific requirements, we introduce modifications by replacing DB with CSDB, which is better tailored for this application.
Additionally, most existing ED models allocate tens of millions of parameters to feature extraction but rely on a single $1\times 1$ convolution layer for feature fusion. This imbalance may lead to underutilization of extracted features, potentially limiting performance. To address this, unlike previous models that typically fuse multiple outputs using a single $1\times 1$ convolution layer, the proposed model employs a fusing block consisting of three convolution layers, including a $3\times 3$ convolution layer, to enhance feature integration.
@@ 3.1 The SDPED Model
The model is trained using the weighted binary cross-entropy (WBCE) loss function, defined as:
$L_{WBCE}(\hat{Y},Y)=-\alpha\sum_{y_{i}\in Y^{+}}\log(\hat{y_{i}})-\lambda(1-\alpha)\sum_{y_{i}\in Y^{-}}\log(1-\hat{y_{i}}),$ (4)
where $\hat{Y}$ represents the model’s prediction, $Y^{+}$ is the set of positive samples (edge pixels in the ground-truth), and $Y^{-}$ is the set of negative samples (non-edge pixels in the ground-truth). The weighting factor $\alpha$ is computed as:
@@
Figure 3: Cascading Skipping Dense Block (CSDB): Each CSDB block consists of three cascaded SDB blocks connected via skip connections. Within a single SDB block, there are five convolution layers. The first four layers output 32 feature maps, while the final layer produces 64 feature maps. The input feature dimensions grow progressively from 64 to 96, 128, 160, and 192 across the first four layers.
3.2 Enhancing Performance on Noiseless Data
Existing models, along with their data augmentation techniques, often fail to yield satisfactory results even on noiseless data. Edge maps inherently contain higher-quality edge information than raw images, and thus predictions on such noiseless data should ideally be sharper and more precise. However, as shown in Fig. 1, previous models tend to produce similar outputs for both noisy and noiseless inputs.
The primary reason for this limitation is that existing models are trained on datasets where label noise is inevitable. Like all supervised learning methods, ED models are highly dependent on the quality of training data. However, obtaining high-quality ED datasets is particularly challenging, as edge annotations are manually labeled and inherently noisy, often containing missing or erroneous labels. This labeling noise can significantly degrade model performance, even when evaluated on high-quality test data.
The most straightforward solution to this problem would be incorporating noiseless data into the training set. However, since human annotations are currently the only feasible method for generating ground-truth labels, obtaining completely noiseless data appears impractical.
Nevertheless, a potential solution arises by directly using ground-truth labels as training samples. ED labels are inherently noiseless when treated as input images with their own ground-truth since the edge map of an edge map—regardless of how it is generated—remains unchanged. Leveraging this property, we integrate noiseless data into the training process by using labels as input images, effectively enhancing dataset quality without requiring additional annotations. With this modified data augmentation strategy, models can generate more accurate and visually appealing predictions, particularly on noiseless inputs, as demonstrated in Fig. 1.

### v5

# High-Precision Edge Detection via Task-Adaptive Texture Handling and Ideal-Prior Guidance

@@ Abstract
Image edge detection (ED) requires specialized architectures, reliable supervision, and rigorous evaluation criteria to ensure accurate localization. In this work, we present a framework for high-precision ED that jointly addresses architectural design, data supervision, and evaluation consistency. We propose SDPED, a compact ED model built upon Cascaded Skipping Density Blocks (CSDB), motivated by a task-adaptive architectural transfer from image super-resolution. By re-engineering texture-oriented structures for ED, SDPED effectively differentiates textures from edges while preserving fine spatial precision. Extensive experiments on four benchmark datasets (BRIND, UDED, MDBD, and BIPED2) demonstrate consistent performance improvements, particularly in Average Precision (AP), with gains of up to 22.5% on MDBD and 11.8% on BIPED2. In addition, we introduce an ideal-prior guidance strategy that incorporates noiseless data into training by treating labels as noise-free samples, providing a practical means to mitigate the subjectivity and noise inherent in human annotations. To enable fair and resolution-independent evaluation, we further adopt a fixed-pixel criterion for assessing localization accuracy. Overall, this work offers a coherent solution for high-precision ED and provides insights applicable to precision-oriented modeling in low-level and soft computing–based vision tasks. Codes can be found on https://github.com/Hao-B-Shu/SDPED.
@@
Edge Detection, Task-Adaptive Architecture, Ideal-Prior Guidance, Noiseless Supervision, Evaluation Benchmark.
@@
[…]
@@
A compact edge detection architecture based on cascaded skipping density blocks.
Task-adaptive architectural transfer from super-resolution to edge detection.
An ideal-prior guidance strategy leveraging noiseless annotations to mitigate label noise.
Resolution-independent evaluation for rigorous and consistent edge detection assessment.
Consistent precision improvements validated across multiple datasets and partitions.
@@ 1 Introduction
Image edge detection (ED) is a fundamental task in computer vision and has been studied for decades, with extensive applications in high-level tasks such as object detection[1] and image segmentation[2]. Early ED methods, such as the Sobel operator[3] and the Canny algorithm[4], relied on gradient-based features. Later, more sophisticated statistical approaches were introduced[5, 6]. In recent years, machine learning has become indispensable in nearly all computer vision tasks, leading to the emergence of numerous learning-based ED algorithms[7, 8, 9, 10, 11, 12]. Currently, convolutional neural network (CNN)-based models achieve state-of-the-art (SOTA) performance in ED tasks [13, 14, 15, 16, 17, 18, 19], while recently, transformer-based approaches have also been explored[20].
The goal of ED is to identify edges in an image, which typically correspond to pixels with significant changes in intensity compared with their neighbors. However, textural areas can also contain rapid intensity changes that do not correspond to edges. Classical ED algorithms such as Canny[4], despite their efficiency, often fail to differentiate between edges and textures due to their inability to adapt to varying data distributions. Consequently, learning-based models, particularly CNNs, have become the dominant approach. Over the past decade, CNN-based ED models have demonstrated remarkable progress, with models such as HED[14], RCF[21], RINDNet[22], BDCN[23], DexiNed[19], and even lightweight models such as TEED[24] achieving $F$-scores comparable to or exceeding human performance under commonly adopted evaluation procedures.
@@
    Insufficient Precision under Strict Constraints: Current ED models achieve high scores only when evaluated with a large error tolerance. For example, the standard error tolerance radius is approximately 4.3 pixels for BSDS500[6], 11.1 pixels for BIPED[25], and even 16.3 pixels for MDBD[26]. However, when stricter evaluation criteria are applied to ensure high-precision localization, the performance of these models drops significantly. This indicates that current architectures may lack sufficient spatial fidelity to distinguish fine edges from surrounding textures, suggesting the need for architectures explicitly designed for high-precision localization.
@@
    Suboptimal Reliability on Noiseless Inputs: , As shown in Fig 1, existing models often produce suboptimal edge maps even when given noiseless input data, namely ground-truth edge maps themselves, which represent deterministic and noise-free ideal priors. Ideally, an ED model should yield nearly perfect predictions on noiseless inputs since they represent the ”ideal prior” of edges. Yet, most models generate results on noiseless inputs that are similar to those on textured images, suffering from blurred or discontinuous predictions. This limitation fundamentally stems from training solely on human-annotated data, where subjectivity and annotation noise are unavoidable.
Figure 1:    Predictions of three previous ED models on the BRIND dataset. $Off_{E}$ denotes models retrained using the official data augmentation and tested on edge maps, while $Our_{E}$ represents models retrained using our proposed data augmentation (detailed in Section 3.2) and tested on edge maps. Similarly, $Off_{O}$ and $Our_{O}$ correspond to models trained with respective augmentation strategies but tested on standard images. Our augmentation method produces sharper predictions on edge maps and potentially enhances performance on standard images.
To address these challenges, we propose a unified high-precision ED framework. We observe a fundamental synergy between image super-resolution (SR) and ED: both tasks rely on carefully handling textures, albeit with different task objectives. Drawing inspiration from this, we perform a task-adaptive architectural transfer by introducing Cascaded Skipping Density Blocks (CSDB). While density blocks are adept at handling textures in SR[27], our investigation shows that their direct application to ED is insufficient for achieving precise localization. However, by re-engineering them into a cascaded skipping structure with task-aware modifications, the architecture can be effectively transferred to meet the stringent precision requirements of ED. Furthermore, we introduce a novel data augmentation strategy that integrates noiseless annotations as ideal prior guidance, which allows the model to learn from deterministic, noise-free samples as ideal guidance, thereby alleviating the performance ceiling imposed by subjective human annotations. To ensure a scientifically rigorous assessment of these improvements, we also advocate for and analyze the necessity of employing stricter evaluation standards to measure true localization precision.
@@
    Task-Adaptive High-Precision Architecture: We introduce Skipping Density Precise Edge Detection (SDPED), a parameter-efficient yet powerful model for high-performance edge detections. By incorporating CSDB—a structure not previously adapted to ED in a task-aware manner—we demonstrate that dense architectural mechanisms from SR can be transferred to ED tasks with several modifications. Notably, our models do not require multi-scale supervision, challenging the widely held belief that such supervision is essential for extracting complex edge features.
@@
    Ideal Prior Guidance via Noiseless Data: We propose a novel data augmentation method that treats training labels as noise-free input samples. As far as we know, this provides the first feasible way to introduce ”perfect knowledge” into ED training, which aims to mitigate the impact of human-annotation noise. This paradigm of label-to-data learning is potentially generalizable to a broad range of supervised tasks in image processing and related domains, especially in tasks where ground-truth labels can only be obtained by human annotations, which unavoidably induces noise.
@@
    Comprehensive Validation and Scientific Evaluation: Through extensive experiments on multiple datasets, we demonstrate that our framework achieves SOTA performance under the strictest benchmark, particularly in Average Precision (AP). Moreover, we also provide an analysis demonstrating that such strict evaluations are essential for fair and high-precision comparison of ED models.
@@ 2.1 Datasets, Label Refinement, and Data Augmentation
Early ED datasets were primarily derived from segmentation tasks, such as BSDS300 and its later extension, BSDS500[28], which remains one of the most widely used benchmarks. BSDS500 consists of 500 RGB images annotated by multiple users and is frequently utilized for both segmentation and boundary detection tasks. Another well-known dataset is NYUD/NYUD2[29], which contains hundreds of thousands of RGB-D images, originally designed for indoor scene analysis. Among these, 1,449 images are annotated with segmentation categories. Other commonly used datasets include PASCAL VOC[30], Microsoft COCO[31], SceneParse150[32], and Cityscape[33]. However, since these datasets were not explicitly designed for ED, their labels often lack the spatial precision required for accurate edge localization.
[…]
A fundamental and persistent challenge across all ED datasets is annotation noise. Since edge labels are based on human annotations, perfect accuracy is unattainable, and inconsistencies inevitably arise both across different annotators and even within the same annotator across multiple attempts. As a result, various studies have focused on refining human-labeled data or enhancing model robustness to noisy annotations[34, 35].
Regardless of the dataset used, data augmentation is widely recognized as essential in ED tasks. Common augmentation techniques include rotation, flipping, cropping, resizing, and gamma transformations. However, augmentation strategies vary across studies. In HED[14], images are rotated at 16 different angles, flipped, and cropped to obtain 32 augmented images per sample. In RIND[22], only four rotations with flips (without cropping) are applied, resulting in eight augmented images per sample. In contrast, Dexi[19] employs a more aggressive augmentation strategy, incorporating 16 rotations, flips, cropping, and three gamma transformations, generating 288 augmented images per sample. While extensive augmentation increases data diversity, its actual impact on edge localization performance remains an open question, and excessive augmentation can significantly increase training costs. Moreover, existing augmentation strategies operate exclusively on human-annotated data and inherently propagate annotation noise. To the best of our knowledge, no existing augmentation strategy incorporates noiseless training data to guide the model toward learning the accurate edge.
@@ 2.2 CNN-based ED Models and Loss Functions
While recent studies have begun exploring transformer-based models[20, 36], CNN-based architectures continue to dominate the ED field. The HED model[14] introduced a multi-scale and multi-level output strategy, where predictions at different levels are fused, and the loss is computed as the average of all supervised outputs. RCF[21] extends this approach by integrating multi-scale outputs with feature fusion across different layers, generating side outputs to enhance edge representation. BDCN[23] applies layer-specific supervision, further improving ED accuracy. PiDiNet[17] introduces pixel-difference convolutions, while RIND[22] predicts four edge types through a three-stage processing pipeline. Dexi[19] leverages extremely dense skip connections to enhance feature propagation.
Most ED models struggle to produce crisp and well-localized edges without additional post-processing techniques such as non-maximal suppression (NMS). This issue may arise from the use of down-sampling operations such as max-pooling, as well as the widespread use of weighted binary cross-entropy (WBCE) as the loss function. However, techniques such as NMS are non-continuous operations and thus cannot be integrated into training pipelines, reducing the value of ED as a reliable and differentiable support cue for high-level vision tasks. Various strategies have been proposed to address this challenge. For instance, [37] suggests improving edge sharpness by aggregating features at larger scales during up-sampling, while [38] explores the effectiveness of Dice loss in ED tasks. More recently, tracing loss[18] has emerged as an alternative and has been adopted in TEED[24], and rank loss[39] decides edges based on ordering pixels. However, these solutions often introduce trade-offs, requiring careful model-specific or dataset-specific hyperparameter tuning, which limits their robustness and general applicability.
@@ 2.3 Evaluation Metrics for ED
Although a variety of evaluation metrics have been explored for ED tasks, such as root mean square error (RMSE), structural similarity index measure (SSIM), and edge ratio (ER)[40], the most-used evaluation framework follows the three key metrics: Optimal Dataset Score (ODS), Optimal Image Score (OIS), and Average Precision (AP). These metrics are derived from the $F_{\beta}$-score, defined as:
@@
[…]
@@ 3 Methodology
This section presents a comprehensive framework for high-precision edge detection (ED). Our approach addresses the limitations of current ED systems through a three-fold synergy: (1) Architectural Design: The proposed SDPED model performs a task-adaptive architectural transfer and modification by re-engineering texture-handling structures from super-resolution (SR) to enhance spatial precision. (2) Learning Paradigm: A novel data augmentation strategy integrates noiseless ideal priors into the training process, breaking the performance ceiling imposed by subjective human annotations. (3) Evaluation Rigor: We analyze the limitations of resolution-dependent error tolerance and discuss fixed-pixel benchmarks as a more rigorous alternative for assessing localization precision.
3.1 The SDPED Model: Task-adaptive Architectural Transfer and Modification
The core challenge of high-precision ED lies in the delicate balance between suppressing complex textural regions and accurately localizing true edges. Since textures lack a formal mathematical definition, designing architectures that can distinguish them from edges remains largely an empirical challenge. However, we observe that image super-resolution (SR), a related low-level vision task, has produced numerous architectures experimentally validated for high-frequency feature handling and texture processing. Therefore, transferring and adapting these structural designs to ED tasks represents a reasonable direction.
Guided by this observation, we propose the SDPED model, as illustrated in Figure 2, with structural details provided in Figure 3. Our model is built upon the Cascaded Skipping Density Block (CSDB), extending the structure of Dense Block (DB) from SR tasks [41, 27]. While the standard DB has demonstrated exceptional performance in texture handling for SR tasks, our experiments indicate that its direct application to ED leads to suboptimal localization performance; therefore, we reformulate it into the CSDB structure with task-adaptive modifications tailored to ED. Furthermore, we address a common imbalance in existing ED models: many architectures allocate millions of parameters to feature extraction but rely on a simplistic $1\times 1$ convolution for feature fusion, which often leads to the underutilization of rich semantic features. To address this, SDPED employs a dedicated, extending fusing block consisting of three convolutional layers (including $3\times 3$) to enhance feature integration and refine the final edge map, allowing the rich information captured by the CSDBs to be more effectively leveraged for the final prediction.
@@ 3.1 The SDPED Model: Task-adaptive Architectural Transfer and Modification
As in most ED models, the model is trained using the weighted binary cross-entropy (WBCE) loss function, defined as:
[…]
where $\hat{Y}$ represents the model’s prediction, $Y^{+}$ is the set of positive samples (edge pixels in the ground-truth), $Y^{-}$ is the set of negative samples (non-edge pixels in the ground-truth), and, $y_{i}$ and $\hat{y}_{i}$ represent the corresponding pixels in the ground-truth and prediction, respectively. The weighting factor $\alpha$ is computed as:
@@
Figure 3:    Cascading Skipping Dense Block (CSDB): Each CSDB block consists of three cascaded SDB blocks connected via skip connections. Within a single SDB block, there are five convolution layers. The first four layers output 32 feature maps, while the final layer produces 64 feature maps. The input feature dimensions grow progressively from 64 to 96, 128, 160, and 192 across the first four layers. In represents the input, and, $O_{1}$, $O_{2}$ and $O_{3}$ represent the outputs of each SDB, respectively.
3.2 Ideal Prior Guidance: Enhancing Performance via Noiseless Data
A critical limitation of existing ED models is their inability to yield perfect results even when the input is noiseless (for example, the ground-truth edge map itself). As demonstrated in Fig. 1, models produce blurred predictions on noiseless data, similar to their performance on noisy (textured) images. We argue that this stems from the human-annotation ceiling: since models are trained exclusively on noisy, human labels, they are unlikely to learn the ideal physical representation of a noise-free edge.
The most direct solution—obtaining perfectly clean, objective labels—is practically impossible due to the nature of manual annotation. However, we propose a novel paradigm: label-to-data augmentation, which allows noiseless training labels to be added to training sets. Observing that in the context of ED, a ground-truth label can be treated as a noiseless input image because the edge map of an edge map is identical to itself, we provide the model with ideal prior guidance by incorporating ground-truth labels directly into the training set as input samples. Namely, we add noiseless data to the training set by viewing labels as training data too, whose annotations are themselves.
This provides a practically feasible way, and might be the only way, to introduce deterministic, noise-free knowledge into the training process without requiring additional human effort. It enables models to learn from perfect data-label pairs on certain samples, thereby enhancing the robustness when facing both noisy and noiseless images. Furthermore, this approach is highly generalizable: it can potentially be extended to a wide range of supervised tasks, particularly those where ground-truth labels are inevitably imperfect. As shown in Figure 1, this ideal prior training allows models to generate visually superior predictions.
3.3 Rethinking Evaluation Benchmarks for High-Precision ED
To assess model performance, the community widely adopts the methodology proposed in [7], which computes ODS, OIS, and AP within an error-tolerance distance, denoted by maxDist. While the metrics themselves are standardized, the selection criteria for maxDist in current research are often inconsistent and lack a rigorous theoretical basis, which poses a significant challenge for high-precision ED tasks.
Currently, maxDist is typically defined as a fixed ratio of the image’s diagonal length, which is inherently problematic because it couples the evaluation standard with image resolution. As resolution increases, the error tolerance is relaxed. For instance, a ratio of 0.0075 translates to a 4.3-pixel tolerance for images of resolution $321\times 481$. However, if one were to concatenate four such images into a single $642\times 962$ frame, the same ratio rule would double the allowable error to 8.6 pixels, despite the underlying structures remaining identical. Consequently, simply enlarging images can lead to inflated evaluation scores without any actual improvement in detection performance. Such loose constraints mask the deficiencies of models in fine localization, reducing the scientific persuasiveness of the evaluation.
Moreover, the inconsistency across different datasets further complicates the comparability of ED models. Even when using the same ratio, such as 0.0075, the tolerance varies from 4.3 pixels in BRIND ($321\times 481$) to 11.1 pixels in BIPED ($720\times 1280$). In the MDBD dataset, the standard is often further relaxed by using a ratio of 0.011, reaching up to 16.3 pixels of error tolerance. This means a prediction that would be penalized as a false positive in one dataset might be accepted as a true positive in another solely due to the resolution-based scaling or relaxed error tolerance, which reduces the persuasiveness of model assessment across datasets.
Therefore, a rethinking of the philosophy on using the evaluated algorithm should be conducted. To ensure a resolution-independent evaluation, we propose setting error tolerance as a fixed pixel value rather than a proportional ratio. Furthermore, we contend that models must be evaluated under a low error-tolerance distance, since a high score under a relaxed threshold but a poor one under a strict one indicates that errors are masked and misaligned predictions are contributing disproportionately to the final score [^1] For example, a prediction missing a single edge pixel should ideally score higher than one missing that pixel while introducing an additional false edge elsewhere. However, a loose error tolerance increases the likelihood of the latter receiving a higher score, as the erroneous pixel may still fall within the large tolerance range.. However, the strictest possible standard—a zero-pixel tolerance—is often unrealistic due to the inherent nature of human labels. As illustrated in Figure 4, edge annotations in current datasets are manually labeled and often exhibit a one-pixel deviation. In physical images, a true edge typically occupies a two-pixel width (spanning the boundary between two regions), yet ground-truth labels are usually marked as a one-pixel-wide path. This subjective choice means labels may vary by up to one pixel horizontally, vertically, or diagonally, depending on the annotator’s preference. Therefore, our analysis suggests that zero-pixel tolerance is overly punitive toward legitimate annotation inconsistencies, while relaxed error tolerance risks accepting erroneous noise. Based on these analyses, we identify the 1-pixel threshold as the strictest standard for evaluation, in practice.